HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

CAVITY RINGDOWN SPECTROSCOPY OF DISSOCIATED NAPHTHALENE IN A RADIOFREQUENCY PLASMA JET

J. LECOMTE, N. SUAS-DAVID, E. DUDÁS, S. PEROT, L. RUTKOWSKI, \underline{R. GEORGES} , IPR, Université de Rennes, UMR6251 CNRS, F-35042, France; C. CHARLES, R. W. BOSWELL, Research School of Physics, Australian National University, Canberra, ACT, Australia; S. KASSI, LIPhy, UMR 5588, Université

Grenoble Alpes, Saint Martin d’Hères, France

Polycyclic aromatic hydrocarbons (PAHs) are abundant organic molecules detected in several objects in the universe, such as molecular clouds in the interstellar medium (ISM)^{1}. Their structure can be modified through plasma-driven processes occurring in the ISM. The present study focuses on the dissociation of naphthalene (C_{10}H_{8}) in a radiofrequency (RF) plasma, probed using cavity ringdown spectroscopy (CRDS) in the near-infrared. Namely, the low-power RF plasma source, called Platypus, is adapted from a small plasma thruster (Pocket Rocket^{2}) designed by the Space Plasma Power and Propulsion laboratory of the ANU. A stable supersonic jet plasma is generated by expanding a mixture of argon and C_{10}H_{8} into a vacuum chamber through a 20 mm long, 4 mm wide slit nozzle^{3}. The jet-cooled fragmented C_{10}H_{8} is finally probed with the ultra-sensitive CRDS technique. We recorded a spectrum from 5950 to 6120 cm^{−1} composed of several hundred transitions originating from many different molecules, radicals, and probably ions^{4}.

1doi:10.1086/184435, L. J. Allamandola et al., The Astrophysical Journal, 290, L25-L28 (1985).

2doi:10.1088/0963-0252/21/2/022002 , C. Charles and R. W. Boswell, Plasma Sources Science and Technology, 21.2, 022002 (2012).

3theses.hal.science/tel-03341145, E. Dudás, Ph.D Thesis, 149-162 (2021).

4doi:10.1021/acs.jpca.9b00100, M. Alliati et al., The Journal of Physical Chemistry, A 123.10, 2107-2113 (2019).

p-number: p185  Submitted on Thu, 15 Jun 2023 14:54:06 +0200